# Efficient Construction of Model Family through Progressive Training Using Model Expansion

### v1

@@ 4.4 Effectiveness of Maximum Learning Rate Adjustment
In LLM pre-training, learning rate selection involves a careful balance between optimization efficiency and training stability. High learning rates can accelerate convergence and potentially improve final model performance (Takase et al., 2023), but this hyperparameter must be carefully adjusted according to the model size. As models grow larger, they become increasingly sensitive to learning rate settings, typically requiring smaller learning rates to maintain training stability (Touvron et al., 2023; Wortsman et al., 2024). For example, in our preliminary experiments, we observed that a learning rate of $1.5\times 10^{-3}$ enabled effective and stable training for a 1B model. However, when we applied this same learning rate to an 8B model, we encountered loss value spikes that ultimately led to training collapse [^7] A visualization of this training instability with the 8B model is provided in the Appendix C..
The progressive training method has the characteristic of starting training from smaller models. Therefore, it is possible to employ high learning rates to enhance performance during small-model training while lowering the learning rate for larger models to stabilize training. In this experiment, we apply this maximum learning rate adjustment to our 2x Chinchilla law setting, linearly decreasing the maximum learning rate from $1.5\times 10^{-3}$ for the 1B model to $3.0\times 10^{-4}$ for the 8B model, in accordance with the increase in model size, while maintaining a cosine decay schedule during training for all models [^8] The learning rates for each model size are described in Appendix D..
Table 3 presents the results of our maximum learning rate adjustment. Progressive consistently outperforms the Independent counterparts, with more substantial improvements than those observed with no learning rate adjustments (Table 2). As illustrated in Figure 2 (Right), our maximum learning rate adjustment leads to more efficient convergence throughout the training process, achieving 31% FLOPs reduction compared to 26% reduction with no learning rate adjustments.
These results demonstrate the effectiveness of our progressive training approach and how appropriate learning rate adjustments can lead to additional performance gains999 Furthermore, we demonstrate in Appendix E that the effectiveness of our approach persists under post-training setups (SFT+DPO), confirming its robustness for practical applications. .
@@ 5 The Consistency across Model Family
In this section, we explore the consistency of model behaviors across different sizes in our progressively trained model family. We first discuss the advantages of maintaining consistent behaviors across the model family (Section 5.1). We then present empirical analyses of consistency through probability distribution, quantified using KL divergence between adjacent model sizes (Section 5.2). Our findings demonstrate that models trained through our progressive training approach exhibit greater consistency compared to independently trained models.
@@ 5.1 The Potential Effectiveness of Consistency across Model Family
Ensuring consistency of behaviors across a model family offers several practical benefits. First, it simplifies deployment flexibility by ensuring that switching between different parameter sizes does not lead to large, sudden shifts in model outputs or user experience. This is particularly valuable when developers need to adapt a model size to varying computational budgets or runtime constraints, as they can easily scale up or down (e.g., from a smaller to a larger model) without retraining users or extensively revalidating system performance (Srivastava et al., 2020; Echterhoff et al., 2024).
Second, this consistency facilitates more efficient incremental improvements or patches across the entire model family. For instance, if developers collect preference data or build a reward model tailored to one member’s outputs (e.g., by annotating the specific texts that this member tends to generate), they inevitably face a distribution mismatch when reusing these artifacts for other models. However, in a consistent family, differences in generation patterns are small enough that the annotated examples remain sufficiently effective for other models (Guo et al., 2024; Zhou et al., 2024a; Tajwar et al., 2024). Likewise, implementations such as safety or content filters calibrated for one model will likely transfer to others with only minimal adjustments (Inan et al., 2023).
Third, consistency between models of different sizes benefits speculative decoding (Leviathan et al., 2023), a technique designed to accelerate inference by letting a smaller model propose “draft” outputs that the larger model then accepts or refines. When the smaller and larger models share similar probability distributions, the larger model is more likely to approve the smaller model’s drafts, reducing the number of rejections and regenerations and thereby improving overall inference speed (Zhou et al., 2024b). Consequently, consistent behavior across a range of model sizes further contributes to more efficient and user-friendly deployments in real-world applications.
@@ 5.2 Probability Distribution Consistency across Model Family
The consistency of the underlying probability distributions across the model family provides deeper insights into how knowledge and behaviors propagate. Probability distribution consistency indicates that models share similar processes when generating text.
@@
To evaluate the probability distribution consistency across our model family, we measured the KL divergence between adjacent model sizes. KL divergence quantifies how one probability distribution differs from another reference distribution, with lower values indicating greater similarity. For this analysis, we used 10,000 examples from the FineWeb-Edu validation dataset. We calculated the KL divergence $D_{\textrm{KL}}(P_{X_{i}}\parallel P_{X_{i+1}})$ between adjacent models by examining their next-token prediction distributions, where $P_{X_{i}}$ represents the probability distribution over the vocabulary given by model of size $X_{i}$101010 The detailed calculation method for KL divergence is described in Appendix F. We conducted this analysis for three different model families: (i) Models trained independently from scratch, (ii) Models built through progressive training, (iii) Models built through progressive training with maximum learning rate adjustment.
@@ Results.
Table 4 presents the KL divergence between adjacent model sizes for our three different model families. We found that, across all model pairs, models built through progressive training demonstrate substantially lower KL divergence compared to independently trained models, indicating higher consistency in their probability distributions. This trend is consistent across all adjacent model pairs (1B→2B, 2B→4B, 4B→8B).
This finding provides evidence that our progressive training approach constructs a more coherent model family with consistent internal behavior. The lower KL divergence values suggest that progressively trained models share similar probability distributions and decision-making processes, offering practical advantages for techniques like speculative decoding where alignment between smaller and larger models is crucial.
@@ 6 Related Work
Offering models in various sizes has become standard practice in language model development to address diverse computational requirements. Major research labs have released model family with multiple parameter sizes: Llama (Touvron et al., 2023; AI@Meta, 2024), Qwen (Yang et al., 2024a), Gemma (GemmaTeam, 2024), and others. Scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022) provide the theoretical foundation for these families, establishing relationships between model size, data, compute, and performance. Recent research highlights the complementary roles of models at different sizes (Wang et al., 2024a); while large models excel in zero/few-shot generalization, smaller models offer advantages in latency-sensitive applications, edge deployment, domain-specific tasks, and privacy-sensitive contexts. This functional differentiation underscores the importance of efficient methods for constructing a coherent model family that maintains consistent capabilities across different parameter sizes. However, the conventional approach of training each model size independently incurs substantial computational costs that scale additively with the number of models in the family. Our work addresses this inefficiency by proposing progressive training that significantly reduces the total computation required to construct a complete, coherent model family.
@@
Model expansion has emerged as an approach to reduce computational costs when training large-scale models. These methods leverage parameters of smaller, pre-trained models to initialize larger ones, accelerating training. The bert2BERT (Chen et al., 2022) adapted function-preserving methods to Transformers, reducing pre-training costs by approximately 45% for BERT models. LEMON (Wang et al., 2024b) refined this approach with an optimized learning rate scheduler for larger models. For scaling LLMs, Du et al. (2024) identified depthwise stacking as particularly effective, achieving 54.6% speedup for 7B-parameter models. However, these methods typically focus on the efficient training of a single, final model rather than the training of an entire model family. For example, many methods demonstrate the effectiveness of the expansion in a single-shot manner, i.e., applying model expansion only once. In addition, these methods do not take the performance of intermediate model(s) into account. In contrast, we propose progressive training and demonstrate that the off-the-shelf model expansion methods can be used to efficiently train a model family. The progressive training can incorporate emerging novel model expansion techniques, further enhancing efficiency.
@@ 7 Conclusion
In this paper, we propose an efficient method for constructing a model family through progressive training, where smaller models are incrementally expanded to larger sizes. Through comprehensive experiments with the model family ranging from 1B to 8B parameters, we demonstrated that our approach significantly reduces the total computational cost required to construct a complete model family by approximately 25% compared to training each model independently. Our method, particularly when combined with maximum learning rate adjustments tailored to each model size, not only matches but exceeds the performance of independently trained models. Furthermore, models built through our progressive approach demonstrate greater consistency across different sizes, as evidenced by lower KL divergence between probability distributions.
This work provides researchers and practitioners with a computationally efficient approach to constructing a high-performing, coherent model family, offering particular value in resource-constrained environments where training multiple models at different scales is desirable.
@@ Appendix D Specific Configurations in the Learning Rate Adjustment Strategy
In our learning rate adjustment strategy (Section 4.4), we decrease the maximum learning rate as the model size increases. We apply maximum learning rates of $1.5\times 10^{-3}$, $1.1\times 10^{-3}$, $7.0\times 10^{-4}$, and $3.0\times 10^{-4}$ for the 1B, 2B, 4B, and 8B models, respectively. This gradual reduction accommodates the increasing sensitivity of larger models to high learning rates.

### v2

@@ 4.4 Effectiveness of Maximum Learning Rate Adjustment
In LLM pre-training, selecting an appropriate learning rate requires a careful balance between optimization efficiency and training stability. Higher learning rates can accelerate convergence and potentially lead to better final model performance (Takase et al., 2023), but they must be tuned carefully, especially with respect to model size. As models scale up, they become increasingly sensitive to the learning rate, typically requiring smaller learning rates to preserve training stability (Touvron et al., 2023; Wortsman et al., 2024). For example, in our preliminary experiments, we observed that a learning rate of $1.5\times 10^{-3}$ enabled effective and stable training for a 1B model. However, applying the same learning rate to an 8B model resulted in the loss value spikes that ultimately led to training collapse [^7] A visualization of this training instability with the 8B model is provided in the Appendix C..
Progressive training method is characterized by starting training from smaller models. Therefore, it is possible to employ high learning rates to enhance performance during small-model training while lowering the learning rate for larger models to stabilize training. In this experiment, we apply this maximum learning rate adjustment under the 2x Chinchilla law setting, linearly decreasing the maximum learning rate from $1.5\times 10^{-3}$ for the 1B model to $3.0\times 10^{-4}$ for the 8B model, in accordance with the increase in model size, while maintaining a cosine decay schedule during training for all models [^8] The learning rates for each model size are described in Appendix D..
Table 3 presents the results of our maximum learning rate adjustment strategy. Progressive consistently outperforms the Independent counterparts, with more substantial improvements than those observed with no learning rate adjustments (Table 2). As illustrated in Figure 2 (Right), our maximum learning rate adjustment facilitated more efficient convergence during training, achieving 31% FLOPs reduction compared to 26% reduction observed without learning rate adjustment. These results demonstrate the effectiveness of our progressive training approach and show that appropriate learning rate adjustments can further enhance performance [^9] Moreover, as shown in Appendix E the benefits of our approach persist under post-training setups (SFT+DPO), confirming its robustness for practical applications. .
@@ 5 The Consistency across Model Family
In this section, we investigate the consistency of model behaviors across different sizes in our progressively trained model family. We first discuss the advantages of maintaining consistent behaviors across the model family (Section 5.1). We then present empirical analyses of consistency through probability distribution and speculative decoding. Our findings show that models trained using the progressive training approach exhibit higher consistency compared to independently trained models.
@@ 5.1 The Potential Effectiveness of Consistency across Model Family
Ensuring consistent behavior across a model family offers several practical benefits. First, it simplifies deployment flexibility by ensuring that switching between different parameter sizes does not lead to large, sudden shifts in model outputs or user experience. This is particularly valuable when developers need to adapt a model size to varying computational budgets or runtime constraints, as they can easily scale up or down (e.g., from a smaller to a larger model) without retraining users or extensively revalidating system performance (Srivastava et al., 2020; Echterhoff et al., 2024).
Second, such consistency enables more efficient incremental improvements or patches across the entire model family. For instance, if developers collect preference data or build a reward model based on the outputs of one model (e.g., by annotating the specific texts that this member tends to generate), they typically face a distribution mismatch when applying those artifacts to other models. In a consistent model family, however, differences in generation patterns are small enough that examples or learned preferences remain effective for other models (Guo et al., 2024; Zhou et al., 2024a; Tajwar et al., 2024). Likewise, implementations such as safety or content filters calibrated for one model will likely transfer to others with only minimal adjustments (Inan et al., 2023).
Third, consistency across models of different sizes benefits speculative decoding (Leviathan et al., 2023), a technique that accelerates inference by allowing a smaller model to generate “draft” outputs, which the larger model then either accepts or refines. When the smaller and larger models produce similar probability distributions, the larger model is more likely to accept the drafts, reducing the frequency of rejections and subsequent regenerations (Zhou et al., 2024b). Consequently, consistent behavior across a range of model sizes further contributes to more efficient and user-friendly deployments in real-world applications.
@@ 5.2 Probability Distribution Consistency across Model Family
The consistency of the underlying probability distributions across the model family offers deeper insights into how knowledge and behaviors are propagated. Such consistency indicates that models rely on similar internal mechanisms when generating text.
@@
To evaluate the probability distribution consistency across our model family, we measured the KL divergence between adjacent model sizes. KL divergence quantifies how one probability distribution differs from another reference distribution, with lower values indicating greater similarity. For this analysis, we used 10,000 examples from the FineWeb-Edu validation dataset. We calculated the KL divergence $D_{\textrm{KL}}(P_{X_{i}}\parallel P_{X_{i+1}})$ between adjacent models by examining their next-token prediction distributions, where $P_{X_{i}}$ represents the probability distribution over the vocabulary given by model of size $X_{i}$ [^10] The detailed calculation method for KL divergence is described in Appendix F. We conducted this analysis for three different model families: (i) Models trained independently from scratch, (ii) Models built through progressive training, (iii) Models built through progressive training with maximum learning rate adjustment.
@@ Results.
Table 4 presents the KL divergence between adjacent model sizes for our three different model families. We found that, across all model pairs, models trained via progressive training demonstrate substantially lower KL divergence compared to independently trained models, indicating higher consistency in their probability distributions. This trend is consistent across all adjacent model pairs (1B→2B, 2B→4B, 4B→8B).
This finding provides evidence that our progressive training produces a more coherent model family with consistent internal behavior. The lower KL divergence suggests that progressively trained models share similar probability distributions and decision-making processes, offering practical advantages for techniques like speculative decoding, where alignment between smaller and larger models is crucial.
5.3 Validating Consistency Benefits through Speculative Decoding
We empirically assess whether this consistency translates into practical benefits for speculative decoding, as discussed in Section 5.1.
Experimental Setup.
In speculative decoding, a smaller “draft” model proposes multiple tokens that a large “generator” model subsequently accepts or rejects. For our experiments, we fixed the 8B model by Progressive training as the generator and paired it with 4B draft models trained using either Progressive or Independent training. Since the effectiveness of speculative decoding depends on how well the draft model’s predictions align with the generator’s distribution, this setup validates whether the distributional consistency yields practical benefits. Each draft model was configured to propose 8 speculative tokens. To evaluate speculative decoding performance, we measured three key metrics on 1,000 prompts sampled from the FineWeb-Edu validation dataset. First, we computed the KL divergence $D_{\textrm{KL}}(P_{X_{\textrm{draft}}}\parallel P_{X_{\textrm{generator}}})$ between the draft and generator models’ output distributions. Second, we tracked the acceptance rate, which is the percentage of draft tokens accepted by the generator model. This metric directly reflects the efficiency of speculative decoding. Finally, we recorded the average generation time across all prompts to quantify the practical speedup.
Results.
Draft Model Generator Model $D_{\textrm{KL}}(P_{\text{draft}}\parallel P_{\text{generator}})$ Acceptance Generation
(4B) (8B) Rate (%) Time (s)
Progressive Progressive 0.3378 93.20 7.02
Independent Progressive 0.4281 87.20 8.24
Table 5: Speculative decoding performance with different draft-generator configurations. Progressive draft models demonstrate lower KL divergence with the generator, resulting in higher acceptance rates and faster generation times.
Table 5 presents the speculative decoding performance with different draft-generator configurations. The results demonstrate the practical advantages of employing a draft model constructed through Progressive Training. The pairing of a 4B Progressive draft model with an 8B Progressive generator model (Prog-Prog) exhibited superior performance, underpinned by greater consistency between the two models. This enhanced consistency is quantified by a lower KL divergence for the Prog-Prog pair, compared to using an independently trained model as a drafter (Inde-Prog). This improved alignment in output distributions directly translated to more effective speculative decoding: the Prog-Prog configuration achieved a higher acceptance rate of 93.20%, as opposed to 87.2% for the Inde-Prog setup. Consequently, this led to a reduction in inference time, with the Prog-Prog pair completing generation approximately 14.8% faster. These findings experimentally demonstrate that the consistency fostered by Progressive Training offers practical benefits. The lower KL divergence between progressively trained models leads to more effective draft proposals in speculative decoding, thereby improving acceptance rates and reducing inference latency.
@@ 6 Related Work
Offering models in multiple sizes has become a standard practice in language model development to address diverse computational requirements. Major research labs have released model families with multiple parameter sizes: Llama (Touvron et al., 2023; AI@Meta, 2024), Qwen (Yang et al., 2024a), Gemma (GemmaTeam, 2024), and others. Scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022) provide the theoretical foundation for these families, establishing relationships between model size, data, compute, and performance. Recent research highlights the complementary roles of models of different sizes (Wang et al., 2024a): while large models excel in zero/few-shot generalization, smaller models offer advantages for latency-sensitive applications, edge deployments, domain-specific tasks, and privacy-sensitive contexts. This functional differentiation underscores the importance of efficient methods for constructing a coherent model family that maintains consistent capabilities across different parameter sizes. However, the conventional approach of training each model size independently incurs substantial computational costs that scale additively with the number of models in the family. Our work addresses this inefficiency by proposing progressive training that significantly reduces the total computation required to construct a complete and coherent model family.
@@
Model expansion has emerged as an approach to reduce computational costs when training large-scale models. These methods leverage parameters of smaller, pre-trained models to initialize larger ones, accelerating training. The bert2BERT (Chen et al., 2022) adapted function-preserving methods to Transformers, reducing pre-training costs by approximately 45% for BERT models. LEMON (Wang et al., 2024b) refined this approach with an optimized learning rate scheduler for larger models. For scaling LLMs, Du et al. (2024) identified depthwise stacking as particularly effective, achieving 54.6% speedup for 7B-parameter models. Alternative approaches combine knowledge distillation with model expansion.  Qin et al. (2022) introduced Knowledge Inheritance (KI), using knowledge distillation during pre-training to achieve approximately 27% computational cost reduction. However, these methods typically focus on the efficient training of a single, final model rather than the training of an entire model family. For example, many methods demonstrate the effectiveness of the expansion in a single-shot manner, i.e., applying model expansion only once. In addition, these methods do not take the performance of intermediate model(s) into account. In contrast, we propose progressive training and demonstrate that the off-the-shelf model expansion methods can be used to efficiently train a model family. The progressive training can incorporate emerging novel model expansion techniques, further enhancing efficiency.
@@ 7 Conclusion
In this paper, we propose an efficient approach for constructing a model family via progressive training, wherein smaller models are incrementally expanded to larger sizes. Through comprehensive experiments on a model family ranging from 1B to 8B parameters, we demonstrated that our method significantly reduces the total computational cost required to construct a complete model family by approximately 25% compared to training each model independently. Our method, particularly when combined with maximum learning rate adjustments tailored to each model size, not only matches but exceeds the performance of independently trained models. Furthermore, models built through our progressive approach demonstrate greater consistency across different sizes, as evidenced by lower KL divergence between probability distributions.
This work offers researchers and practitioners a computationally efficient approach to constructing a high-performing, coherent model family, offering particular value in resource-constrained environments where training multiple models at different scales is desirable.
Acknowledgements
This work was supported by the JST Moonshot R&D Grant Number JPMJMS2011-35 (fundamental research).
@@ Appendix D Specific Configurations in the Learning Rate Adjustment Strategy
In our learning rate adjustment strategy (Section 4.4), we decrease the maximum learning rate as the model size increases. We apply maximum learning rates of $5\times 10^{-3}, 1.1$$1\times 10^{-3}$, $7.0\times 10^{-4}$, and $3.0\times 10^{-4}$ for the 1B, 2B, 4B, and 8B models, respectively. This gradual reduction accommodates the increasing sensitivity of larger models to high learning rates.